# Multiscale Growth Kinetics of Model Biomolecular Condensates Under Passive and Active Conditions

### v1

@@ 2.3 Network-like local structure under active condition
In the subsequent series of experiments, we examine the growth kinetics of condensates under active condition. The serine residues of the RRASLRRASL peptide are initially phosphorylated, which results in a neutral net charge. As a consequence, mixing phosphorylated peptides and RNA does not give rise to phase separation. Phase separation is instead triggered by the action of lambda phosphatase (LPP), which cleaves the phosphate groups from phosphorylated peptides. Figure 5a plots scattering curves collected under the same concentrations as in Fig. 2a in the presence of 800 U.mL-1 LPP and using initially phosphorylated peptide (see Fig. LABEL:fig:SI-tr-saxs-active for other peptide concentrations). The condensate formation is much slower than under passive condition because phase separation is now limited by the LPP enzymatic activity. Secondly, the scattering curves in the intermediate $q$-values, i.e., between 10-2 and 10-1 nm-1, differ markedly. The inset of Figure 5a presents a comparison between two scattering curves with similar $I_{0}$ values. For actively-formed condensates, $I(q)$ decays at smaller $q$-values indicating large condensate sizes, but the slope is moderate ($I\sim q^{-2.4}$), which suggests a local mass fractal structure akin to that found in gels. Therefore, instead of fitting the scattering curves with a model of polydisperse spheres as earlier, we apply a model of polydisperse mass fractal clusters with a stretched exponential cutoff function 37, 36. The form factor of the clusters is expressed by:
@@
where $D$ stands for the fractal dimension and $h(r;\xi)$ is a cutoff function setting the cluster size $\xi$. The denominator ensures the normalization condition, namely, $F(q\rightarrow 0;\xi)=1$. We choose a generic form for the cutoff function, i.e., $h(r;\xi)=e^{-(r/\xi)^{\beta}}$ with $\beta=3$ as it fits satisfactorily the scattering curves at late times. As before (Eq. 1), the scattering intensity is computed as a convolution product of the form factor and a lognormal size distribution defined by $p(\xi;\langle\xi(t)\rangle,\Delta\xi(t))$, so that
$I(t,q)\sim\displaystyle\int_{0}^{+\infty}F(q;\xi)p(\xi;\langle\xi(t)\rangle,\Delta\xi(t))\xi^{2D(t)}d\xi$ (3)
with a prefactor ensuring that $I(t,q\rightarrow 0)=I_{0}(t)$. In practice, the scattering curves are fitted over the region $q<\text{0.1 nm\textsuperscript{-1}}$, which corresponds to length scales larger than a single RNA chain. The blue dashed lines of Figure 5a represents fits to the experimental data (black solid lines), providing estimates of $\langle\xi(t)\rangle$, $\Delta\xi(t)$ and $D(t)$ for each condition. Figure 5b depicts the time evolution of $\langle\xi\rangle$ for three peptide concentrations. We observe mostly two steps: (i) Condensates with size varying around 100 nm are formed from the first minutes after mixing. These condensates display a network-like local structure as evidenced by a mass fractal dimension $D$ below 2 (Fig. 5c). After some period of maturation during which $D$ increases steadily while $\langle\xi\rangle$ remains unchanged, the condensates abruptly start to increase in size, with $\langle\xi\rangle$ seemingly converging toward the classical coarsening behavior scaling as $t^{1/3}$ (black dashed line in Fig. 5b). During the latter stage, $D$ remains stable above 2.5, which indicates that the condensates are already relatively dense and compact. Notice that at 600 µM peptide concentration, no growth is observed for at least the first 2,000 s. Increasing the LPP concentration from 800 U.mL-1 to 1,200 U.mL-1 (gray discs in Figs. 5b and c) shortens the maturation stage consistently with higher enzymatic activity. Finally, it is worth mentioning that although the condensate polydispersity $\Delta\xi/\langle\xi\rangle$ varies significantly during the maturation stage, it tends to a nearly constant value during the growth stage (Fig. LABEL:fig:SI-polydispersity_active), consistently with the self-similar size distribution encountered in coarsening mechanisms.
@@
Next, we investigate the evolution of actively-formed condensates over extended length and time scales by confocal microscopy. Figures 6a-d depicts confocal microscopy images acquired at various times for a peptide concentration of 1,000 µM. Figure 6e represents the time evolution of the mean radius $\langle R\rangle$ and the standard deviation calculated from image analysis. $\langle R\rangle$ increases until $\sim$60 min and remains constant afterwards. Once again, the condensates attain a steady state due to a self-regulation driven by coulombic repulsion. Note that the radii deduced by confocal microscopy (Fig. 6e) are underestimated compared to those obtained by X-ray scattering (Fig. 5b) because the sizes are near the resolution limit of the confocal setup. At a lower peptide concentration of 600 µM, confocal microscopy reveals that 2 h are necessary for the condensate mean radius to reach 1 µm (Fig. LABEL:fig:SI-confocal_active_600_uM). Interestingly, the normalized size distributions, $R/\langle R\rangle$, computed at various times, are mostly identical, which supports the self-similarity hypothesized earlier with TR-USAXS data. Additionally, these distributions are slightly asymmetric, resembling the lognormal distribution expected for a coalescence mechanism 19. This is further illustrated by Figs. 6g and h showing three coalescence events occurring over 4 min.
@@ 2.4 Actively-formed condensates exhibit faster nanoscale dynamics
Intrigued by the difference of local structures uncovered by X-ray scattering, we carry out fluorescence recovery after photobleaching (FRAP) measurements to estimate the diffusivity of TAMRA-peptide within both passively- and actively-formed condensates in the steady state. Figures 7a-d demonstrate the liquid-like nature of condensates with a fluorescence recovery obtained within just a few seconds. By integrating the fluorescence intensity over condensates and fitting the time evolution with an exponential decay function, the diffusivity of TAMRA-peptide, $D_{\mathrm{c}}$, is estimated (Figs. 7e and f). TAMRA-peptide diffuses twice as fast within actively-formed condensates (0.024 µm2.s-1) as within passively-formed condensates (0.012 µm2.s-1). This finding agrees well with the loosely-packed, network-like local structure measured for actively-formed condensates, which confers a facilitated translational mobility to peptides.
Figure 7: FRAP measurements on passively- and actively-formed condensates in the steady state. Peptide concentration is 1,000 µM including 0.5% of TAMRA-peptide. For active condition, 800 U.mL-1 of LPP is used. (a-d) Confocal microscopy images of an actively-formed condensate at various times after photobleaching. Normalized fluorescence intensity versus time for passively- (e) and actively-formed (f) condensates. A dozen condensates are used for each condition. Error bars are standard deviations. The diffusivities $D_{\mathrm{c}}$ are deduced by fitting the data with an exponential decay function (red line).
@@ 3 Conclusion
For passively-formed condensates, peptides bear a positive net charge and bind on RNA in less than a millisecond after mixing. Peptide-decorated RNA then serve as subunits and remarkably behave as an immiscible fluid in a solvent. After nucleation, spherically-shaped, compact condensates steadily grow following the classical coarsening scaling law in $\langle R\rangle=(Kt)^{1/3}$. The prefactor $K$, supported by confocal microscopy images and MD simulations, suggest a Brownian motion-induced coalescence mechanism. Eventually, coalescence is impaired by coulombic repulsion, leading to a steady state.
@@
As for actively-formed condensates, phosphorylated peptides bear a zero net charge and bind on RNA upon phosphate cleavage by LPP. Resultantly, RNA chains are slowly crosslinked by dephosphorylated peptides making up loosely-packed condensates with a network-like local structure. Recall that a phosphorylated peptide bears two serines, each of which carries one phosphate group cleavable by LPP independently of the other. Under the action of LPP, peptides can be dephosphorylated (two phosphate groups cleaved) or partially phosphorylated (only one phosphate group cleaved). Dephosphorylated peptides are however required to trigger phase separation 34. As dephosphorylation proceeds, more RNA chains are uptaken by condensates whin in turn become more compact. After some maturation period, perhaps once dephosphorylated peptides are in sufficient amount, condensate size abruptly increases driven by coalescence, until the steady state is reached.
@@
Elucidating the nonequilibrium dynamics of biomolecular condensates is essential for understanding the functions and the pathological significance of membraneless organelles in the cell, as well as to design synthetic liquid-like carriers for intracellular delivery of therapeutic mRNA for instance 38. Even though the phase diagram of multicomponent fluid mixtures is inherently complex, particularly in the presence of charged polymers 39, 40, 41, the growth kinetics of our passively-formed condensates obey the ordinary droplet coarsening law ($\langle R\rangle\sim t^{1/3}$) for simple binary mixture. Introducing enzymatic activity delays the onset of the thus reaction-limited growth, but more importantly, kinetic effects modulate the local structure of condensates. A loosely-packed content may be crucial for preventing dynamical arrest that can occur with the highly charged biomolecules found in the cell nucleus 42. In that sense, enzymatic activity may not only regulate the formation and dissolution of biomolecular condensates, but also the viscoelastic properties that govern nanoscale dynamics 43, 42.
@@
A recent study has reported that kinetics govern the large-scale spatial organization of phase-separated DNA droplets 44. In particular, fast condensation was found to promote the formation of hyperuniform structures. Although our passively-formed condensates grow on a faster time scale, we find no clear evidence of hyperuniformity, neither on scattering curves nor on confocal microscopy images (Fig. LABEL:fig:SI-structure_factor_passive_active). Therefore, the other factors that govern the large-scale spatial organization of phase-separating fluids are yet to be identified.
@@
At last, a number of condensates, notably transcriptional condensates, are immersed in a crowded environment that can be modeled as heterogeneous elastic medium. Recent theoretical advances have uncovered a rich diversity of condensate growth kinetics governed by local environmental stiffness 45. Investigating these effects experimentally using model condensates will be an important direction for future research.
@@
We thank Laetitia Gargowitsch for polyU characterizations, Raphaël Maire for his help on hyperuniformity and Antoine Maury for confocal microscopy observations. This work is supported by the Agence Nationale de la Recherche (contract ANR-21-CE30-0001-02). G.T. acknowledges the ESRF for allocating beamtime.
@@
Experimental methods, characterizations of polyU and peptides, supplementary figures with MD simulations, TR-USAXS data and confocal microscopy images.

### v2

@@ 2.3 Network-like local structure under active condition
In the subsequent series of experiments, we examine the growth kinetics of condensates under active condition. The serine residues of the RRASLRRASL peptide are initially phosphorylated, which results in a neutral net charge. As a consequence, mixing phosphorylated peptides and RNA does not give rise to phase separation. Phase separation is instead triggered by the action of lambda phosphatase (LPP), which cleaves the phosphate groups from phosphorylated peptides. Figure 5a plots scattering curves collected under the same concentrations as in Fig. 2a in the presence of 800 U.mL-1 LPP and using initially double phosphorylated peptide (see Fig. LABEL:fig:SI-tr-saxs-active for other peptide concentrations). The condensate formation is much slower than under passive condition because phase separation is now limited by the LPP enzymatic activity. Secondly, the scattering curves in the intermediate $q$-values, i.e., between 10-2 and 10-1 nm-1, differ markedly. The inset of Figure 5a presents a comparison between two scattering curves with similar $I_{0}$ values. For actively-formed condensates, the decay of $I(q)$ at small and medium $q$-values is moderate ($I\sim q^{-2.4}$), which suggests a local mass fractal structure akin to that found in gels, while for passively-formed condensates, the decay slope is more consistent with a surface fractal structure. Therefore, instead of fitting the scattering curves with a model of polydisperse spheres as earlier, we apply a model of polydisperse mass fractal clusters with a compressed exponential cutoff function 38, 36. The form factor of the clusters is expressed by:
@@
where $D$ stands for the mass fractal dimension, and $h(r;\xi)$ is a cutoff function that describes the perimeter of the cluster of size $\xi$ and which must decay faster than any power law. The denominator ensures the normalization condition, namely, $F(q\rightarrow 0;\xi)=1$. We choose a generic exponential form for the cutoff function 36, i.e., $h(r;\xi)=e^{-(r/\xi)^{\beta}}$ with $\beta=3$ as it fits satisfactorily the scattering curves at late times. As before (Eq. 1), the scattering intensity is computed as a convolution product of the form factor and a lognormal size distribution defined by $p(\xi;\langle\xi(t)\rangle,\Delta\xi(t))$, so that
$I(t,q)=\frac{I_{0}(t)}{\langle\xi^{2D(t)}(t)\rangle}\displaystyle\int_{0}^{+\infty}F(q;\xi)p(\xi;\langle\xi(t)\rangle,\Delta\xi(t))\xi^{2D(t)}d\xi+C_{\mathrm{bkg}}$ (3)
In practice, the scattering curves are fitted over the region $q<\text{0.1 nm\textsuperscript{-1}}$, which corresponds to lengthscales above the size of a single RNA chain. The blue dashed lines of Fig. 5a represents fits to the experimental data (black solid lines), providing estimates of $I_{0}(t)$, $\langle\xi(t)\rangle$, $\Delta\xi(t)$ and $D(t)$ for each condition. Figure 5b depicts the time evolution of $\langle\xi\rangle$ for three peptide concentrations. We observe mostly two steps: (i) Condensates with size varying around 100 nm are formed from the first minutes after mixing. These condensates display a network-like local structure as evidenced by a mass fractal dimension $D$ well below 3 (Fig. 5c). After some period of maturation during which $D$ increases steadily while $\langle\xi\rangle$ remains unchanged, the condensates abruptly start to increase in size, with $\langle\xi\rangle$ seemingly converging toward the classical coarsening behavior scaling as $t^{1/3}$ (black dashed line in Fig. 5b). During the latter stage, $D$ remains stable above 2.5, which indicates that the condensates are fairly compact. Notice that at 600 µM peptide concentration, no growth is observed for at least the first 2,000 s. Increasing the LPP concentration from 800 U.mL-1 to 1,200 U.mL-1 (gray discs in Figs. 5b and c) shortens the maturation stage consistently with higher enzymatic activity. Finally, it is worth mentioning that although the condensate polydispersity $\Delta\xi/\langle\xi\rangle$ varies significantly during the maturation stage, it tends to a nearly constant value during the growth stage (Fig. LABEL:fig:SI-polydispersity_active), consistently with the self-similar size distribution encountered classically in coarsening mechanisms.
@@
MD simulations that explicitly account for enzymatic activity on phosphorylated peptides are not computationally feasible due to the inherently slow timescales of the reactions involved. Instead, we model cluster formation in systems containing distinct fixed peptide compositions, thereby mimicking different stages of phosphatase-driven dephosphorylation. Figure 6a presents the composition of the six simulated systems, ranging from S1, which predominantly contains double phosphorylated peptides, to WT, which consists solely of non-phosphorylated peptides as set under passive conditions. The exact peptide compositions are detailed in Table LABEL:tab:SI-composition_condensate_active. As shown in Fig. 6b, decreasing the phosphorylation level enhances the ability of clusters to incorporate RNA chains within the same simulation timeframe. This behavior is expected as coacervation relies on positively charged peptides to compensate for the negative charges of RNA. More intriguingly, the final clusters display increasingly loose and heterogeneous morphologies when a substantial fraction of double or mono phosphorylated peptides is present, in contrast to the compact aggregates formed by non-phosphorylated peptides (Fig. 6c). Figures 6d and e quantitatively support this observation by showing that both the radius of gyration and the asphericity of the final clusters increase with peptide phosphorylation level, regardless of the number of incorporated RNA chains. In summary, higher phosphorylation levels lead to less compact and more extended cluster structures, in agreement with experimental measurements.
Figure 6: Coarse-grained MD simulations of condensate growth mimicking active conditions. (a) Peptide compositions of the six simulated systems. (b) Time evolution of the number of RNA chains incorporated into the largest cluster for each system. Shaded areas represent standard deviations over three simulations. (c) Representative snapshots of clusters for the WT and S3 systems. (d) Radius of gyration ($R_{gyr}$) of clusters containing one, two, or three RNA chains for different peptide compositions. (e) Asphericity of clusters containing one, two, or three RNA chains for different peptide compositions.
Next, we investigate the evolution of actively-formed condensates over extended length- and timescales by confocal microscopy. Figures 7a-d depicts confocal microscopy images acquired at various times for a peptide concentration of 1,000 µM. Figure 7e represents the time evolution of the mean radius $\langle R\rangle$ and the standard deviation calculated from image analysis. $\langle R\rangle$ increases until 60 min and remains constant afterwards. Once again, the condensates attain a steady state due to self-regulation driven by Coulomb repulsion. Note that the radii deduced by confocal microscopy (Fig. 7e) are underestimated compared to those obtained by X-ray scattering (Fig. 5b) because the sizes are near the resolution limit of the confocal setup. At a lower peptide concentration of 600 µM, confocal microscopy reveals that 2 h are necessary for the condensate mean radius to reach 1 µm (Fig. LABEL:fig:SI-confocal_active_600_uM). Interestingly, the normalized size distributions, $R/\langle R\rangle$, computed at various times, are mostly superimposed, which supports the self-similarity hypothesized earlier with TR-USAXS data. Additionally, these distributions are slightly asymmetric, resembling the lognormal distribution expected for a coalescence mechanism 19. This is further illustrated by Figs. 7g and h showing three coalescence events occurring over 4 min.
@@ 2.4 Actively-formed condensates exhibit faster nanoscale dynamics
Intrigued by the difference of local structures uncovered by X-ray scattering, we carry out fluorescence recovery after photobleaching (FRAP) measurements to estimate the diffusivity of TAMRA-peptide within both passively- and actively-formed condensates in steady state. Figures 8a-d demonstrate the liquid-like nature of condensates with a fluorescence recovery obtained within just a few seconds. By integrating the fluorescence intensity over condensates and fitting the time evolution with an exponential decay function, the diffusivity of TAMRA-peptide, $D_{\mathrm{c}}$, is estimated (Figs. 8e and f). TAMRA-peptide diffuses twice as fast within actively-formed condensates (0.024 µm2.s-1) as within passively-formed condensates (0.012 µm2.s-1). This finding agrees well with the loosely-packed, network-like local structure measured for actively-formed condensates, which confers a facilitated translational mobility to peptides.
Figure 8: FRAP measurements on passively- and actively-formed condensates in steady state. Peptide concentration is 1,000 µM including 0.5% of TAMRA-peptide. For active condition, 800 U.mL-1 of LPP is used. (a-d) Confocal microscopy images of an actively-formed condensate at various times after photobleaching. Normalized fluorescence intensity versus time for passively- (e) and actively-formed (f) condensates. A dozen condensates are used for each condition. Error bars are standard deviations. The diffusivities $D_{\mathrm{c}}$ are deduced by fitting the data with an exponential decay function (red line).
@@ 3 Conclusion
For passively-formed condensates, peptides bear a positive net charge and bind on RNA in less than a millisecond after mixing. Peptide-decorated RNA then serve as robust subunits and behave as an immiscible fluid in a solvent. After nucleation, spherical condensates steadily grow following the classical coarsening scaling law in $\langle R\rangle=(Kt)^{1/3}$. The prefactor $K$, as well as confocal microscopy images and MD simulations suggest a Brownian motion-induced coalescence mechanism. Eventually, coalescence is impaired by Coulomb repulsion, leading to steady state. Coalescence is preferred over ripening likely because transferring large peptide-decorated RNA from small to large condensates would be too costly in terms of free energy.
@@
As for actively-formed condensates, double phosphorylated peptides bear a zero net charge and bind on RNA upon phosphate cleavage by LPP. Resultantly, RNA chains are slowly crosslinked by dephosphorylated peptides making up loosely-packed condensates with a network-like local structure. Recall that a double phosphorylated peptide bears two serines, each of which carries one phosphate group cleavable by LPP independently of the other. Under the action of LPP, peptides can be non-phosphorylated (two phosphate groups cleaved) or mono phosphorylated (only one phosphate group cleaved). Non-phosphorylated peptides are however required to trigger phase separation 34. As dephosphorylation proceeds, more RNA chains are incorporated into condensates which concommitantly become more compact. After some maturation period, perhaps once non-phosphorylated peptides are supersaturated, condensate size abruptly increases upon coalescence, until steady state is reached.
@@
Elucidating the nonequilibrium dynamics of biomolecular condensates is essential for understanding the functions and the pathological significance of membraneless organelles in the cell, as well as to design synthetic liquid-like carriers for intracellular delivery of therapeutic mRNA for instance 39. Even though the phase diagram of multicomponent fluid mixtures is inherently complex, particularly in the presence of charged polymers 40, 41, 42, the growth kinetics of our passively-formed condensates obey the yet-simple droplet coarsening law ($\langle R\rangle\sim t^{1/3}$) for binary mixtures. Introducing enzymatic activity delays the onset of the reaction-limited growth, but more importantly, kinetic effects modulate the local structure of condensates. A loosely-packed content may be crucial for preventing dynamical arrest that can occur with the highly charged biomolecules found in the cell nucleus 43. In that sense, enzymatic activity may not only regulate the formation and dissolution of biomolecular condensates, but also the viscoelastic properties that govern nanoscale dynamics 44, 43.
@@
A recent study has reported that kinetics govern the large-scale spatial organization of phase-separated DNA droplets 45. In particular, fast condensation was found to promote the formation of hyperuniform structures. Although our passively-formed condensates grow on a faster timescale, we find no clear evidence of hyperuniformity, neither on scattering curves nor on confocal microscopy images (Fig. LABEL:fig:SI-structure_factor_passive_active). Therefore, the other factors that govern the large-scale spatial organization of phase-separating fluids are yet to be identified.
@@
At last, a number of condensates, notably transcriptional condensates, are immersed in a crowded environment that can be modeled as heterogeneous elastic medium. Recent theoretical advances have uncovered a rich diversity of condensate growth kinetics governed by local environmental stiffness 46. Investigating these effects experimentally using model condensates will be an important direction for future research.
@@
We thank Laetitia Gargowitsch for polyU characterizations, Raphaël Maire for his help on hyperuniformity and Antoine Maury for confocal microscopy observations. We have also benefited from fruitful discussions with Pau Bernado, Christine Doucet and Aurélie Fournet. This work is supported by the Agence Nationale de la Recherche (contract ANR-21-CE30-0001). G.T. acknowledges the ESRF for allocating beamtime. E.R.S is grateful to FAPESP for a visiting research scholarship at Université Paris-Saclay (#23/018385-8).
@@
Experimental methods, characterizations of RNA and peptides, supplementary figures with MD simulations, TR-USAXS data and confocal microscopy images.